# disposition

### v2

@@ 4 Nonlinear Cone Separation Conditions
The key tool to prove our main Bishop-Phelps type scalarization results in vector optimization in Section 5.3 will be the separation of two (not necessarily convex) cones by using Bishop-Phelps type separating cones / separating (seminorm-linear) functions. In this section we present all the necessary basics for nonlinear cone separation.
@@
$(\operatorname{cl}\,S_{A}^{0}(\psi))\cap(\operatorname{cl}\,S_{-K}(\psi))=\emptyset,$ (4.1)
$A\cap(\operatorname{cl}(\operatorname{conv}(-K))=\{0\},$ (4.2)
$0\notin\operatorname{cl}\,S_{-K}(\psi).$ (4.3)
The following proposition will later play an important role in proving a Bishop-Phelps type scalarization result (namely Corollary 5.28) for vector optimization problems in real locally convex spaces.
Proposition 4.1.
Assume that (4.2) and (4.3) are valid, and one of the following statements is satisfied:
@@
    $\operatorname{cl}\,S_{-K}(\psi)$ =$=\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}$ (4.4)
@@
    $B_{A}\text{ is compact},$ (4.5)
    $S_{A}^{0}\text{ is closed},$ (4.6)
    $A\text{ is closed and }\operatorname{cl}\,B_{A}\text{ is compact},$ (4.7)
    $A\text{ is closed and }\operatorname{conv}(\operatorname{cl}(B_{A})\cup\{0\})\text{ is closed}.$ (4.8)
Then, (4.1) is valid.
Proof 4.2.
a) Assuming $A$ is closed and convex, by (4.2) we get $\operatorname{cl}\,S_{A}^{0}\cap\operatorname{cl}\,S_{-K}\subseteq A\cap\operatorname{cl}(\operatorname{conv}(-K))=\{0\}$, and so by (4.3) we conclude (4.1).
@@
b) Obviously, (4.2) and (4.3) imply $A\cap\operatorname{cl}\,S_{-K}=\emptyset$. Since there is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ with (4.4), we get that $x^{*}(x)<\alpha$ for all $x\in A\cap(\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi})$. Note that $B_{A}\cup\{0\}\subseteq A\cap\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}$, and if $A$ is closed, then $\operatorname{cl}\,B_{A}\cup\{0\}\subseteq A\cap\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}$. Consider cases:
@@
Case 1: (4.5) is valid. By the continuity of $x^{*}$ and the compactness of $B_{A}\cup\{0\}$, we get some $\beta\in(0,\alpha)$ such that $B_{A}\cup\{0\}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\leq\beta\}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$. Hence, $\operatorname{cl}\,S_{A}^{0}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\leq\beta\}$.
Case 2: (4.6) is valid. Then, $\operatorname{cl}\,S_{A}^{0}=S_{A}^{0}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$.
@@
Case 3: (4.7) is valid. By the continuity of $x^{*}$ and the compactness of $\operatorname{cl}\,B_{A}\cup\{0\}$ we get some $\beta\in(0,\alpha)$ such that $\operatorname{cl}\,B_{A}\cup\{0\}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\leq\beta\}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$. Thus, $\operatorname{cl}\,S_{A}^{0}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\leq\beta\}$.
@@
Case 4: (4.8) is valid. Then, $\operatorname{cl}\,S_{A}^{0}\subseteq\operatorname{cl}(\operatorname{conv}(\operatorname{cl}(B_{A})\cup\{0\}))=\operatorname{conv}(\operatorname{cl}(B_{A})\cup\{0\})\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$.
Recalling that (4.4) is valid, we conclude (4.1).
Remark 4.3.
@@
The result in Proposition 4.1 (regarding assumption b)) is a generalization of a result by Kasimbeyli [45, Lem. 3], where the author studied the Euclidean space $Y=\operatorname{\mathbb{R}}^{n}$, a closed cone $A\subseteq Y$ and a closed convex cone $K\subseteq Y$ with $A\cap(-K)=\{0\}$. Moreover, it is assumed that there is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ such that (4.4) is valid (hence $0\notin\operatorname{cl}\,S_{-K}$ is valid). The conclusion in [45, Lem. 3] is given as the condition (4.1) with $\operatorname{bd}\,A$ in the role of $A$.
@@
The condition (4.4) involved in b) of Proposition 4.1 can be fulfilled by certain Bishop-Phelps type cones, as the following lemma shows.
Lemma 4.4.
@@
Case 2: Assume that $\psi(d)<1$. Put $\gamma:=x^{*}(d)\,(\geq\alpha)$. By our assumption (4.9), there is $e\in Y\setminus\{0\}$ with $\psi(e)=1$ and $x^{*}(e)=0$. Then, $x^{*}(d\pm te)=x^{*}(d)\pm tx^{*}(e)=\gamma$ for all $t\geq 0$, and $\psi(d\pm te)\geq|\,\psi(d)-\psi(te)\,|=|\,\psi(d)-|t|\,|\to\infty$ for $t\to\infty$. Taking into account the continuity of $t\mapsto h_{+}(t):=\psi(d+te)$ and $t\mapsto h_{-}(t):=\psi(d-te)$, as well as the conditions $h_{+}(0)=h_{-}(0)=\psi(d)<1$, $\lim_{t\to\infty}\,h_{+}(t)=\lim_{t\to\infty}\,h_{-}(t)=\infty$, by the well-known intermediate value theorem there are $t_{+},t_{-}\in\operatorname{\mathbb{P}}$ with $h_{+}(t_{+})=h_{-}(t_{-})=1$. Since $x^{*}(d+t_{+}e)=x^{*}(d-t_{-}e)=\gamma\geq\alpha$ we infer $d-t_{-}e\in B_{C}$ and $d+t_{+}e\in B_{C}$, hence $d\in\operatorname{conv}(\{d-t_{-}e,\,d+t_{+}e\})\subseteq S_{C}.$
Of course, $S_{C}=D$ and the closedness of $D$ imply $\operatorname{cl}\,S_{C}=S_{C}$.
Remark 4.6.
Kasimbeyli and Kasimbeyli [44, Th. 4.5] stated a similar result as in Lemma 4.4 in a reflexive Banach space setting (with $\psi(\cdot):=||\cdot||$), but note that the assumption given in (4.9), which in this setting means that $Y$ has dimension two or higher, was not considered in [44, Th. 4.5]. In the one dimensional space ($Y:=\mathbb{R}$, ||$\cdot$|| ⋅ |), the condition (4.9) fails and the conclusion of Lemma 4.4 may fail. To see this, consider $Y:=\mathbb{R}$, $x^{*}:=1$ and $\alpha\in(0,1]$. Then, $C=C(x^{*},\alpha)=\mathbb{R}_{+}$ and $B_{C}(|\cdot|)=\{1\}$ as well as $\operatorname{cl}\,S_{C}(|\cdot|)=\{1\}$ while $\{x\in\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}=[\alpha,1]$.
To prove some more specific results in the normed space setting in the upcoming Proposition 4.11 and Corollary 5.30 (see Section 5.3), we need additional lemmata.
Lemma 4.7.
@@
$\Omega_{2}\subseteq[0,t]\cdot\Omega_{1}.$ (4.10)
@@
Define the function $h:\operatorname{\mathbb{R}}\times Y\to Y$ by $h(s,\omega):=s\cdot\omega$ for all $(s,\omega)\in\operatorname{\mathbb{R}}\times Y$. Since $[0,t]\times\Omega_{1}$ is (weakly) compact in ×$\operatorname{\mathbb{R}}\times Y$, and the function $h$ is (weakly) continuous by the topological vector space properties, we get that $[0,t]\cdot\Omega_{1}$ is (weakly) compact. From (4.10) we derive that the (weakly) closed subset $\Omega_{2}$ of $[0,t]\cdot\Omega_{1}$ is (weakly) compact as well.
Lemma 4.9.
@@
Proof 4.10.
Consider any $||\cdot||^{\prime}\sim||\cdot||$.
@@
$1^{\circ}$. Obviously, $B_{C}(||\cdot||^{\prime})\subseteq C=\operatorname{\mathbb{R}}_{+}\cdot B_{C}(||\cdot||)$. Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $B_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, which shows that $B_{C}(||\cdot||^{\prime})$ is bounded. Moreover, according to [23, Lemma 2.1.43 (iii)], for the bounded and closed norm-base $B_{C}(||\cdot||)$ with $0\notin B_{C}(||\cdot||)$ we have $C=\mathbb{R}_{+}\cdot B_{C}(||\cdot||)$ is closed, hence $B_{C}(||\cdot||^{\prime})$ is closed as well. Then, the compactness of $B_{C}(||\cdot||^{\prime})$ follows from Lemma 4.7 applied for the compact set $\Omega_{1}:=B_{C}(||\cdot||)$ and the closed, bounded set $\Omega_{2}:=B_{C}(||\cdot||^{\prime})$. Note that $0\notin B_{C}(||\cdot||)=\Omega_{1}$.
@@
$2^{\circ}$. Since $C=\operatorname{\mathbb{R}}_{+}\cdot B_{C}(||\cdot||)\subseteq\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||)\subseteq\operatorname{cl}_{w}\,C$, we have $\operatorname{cl}_{w}\,C=\operatorname{cl}_{w}(\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||))$. Taking into account that $\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ is a weakly closed, bounded set with $0\notin\operatorname{cl}_{w}\,B_{C}(||\cdot||)$, applying [23, Lemma 2.1.43 (iii)] we conclude that $\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ is weakly closed (remark that the weak topology of the normed space $Y$ is Hausdorff). Consequently, $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})\subseteq\operatorname{cl}_{w}\,C=\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||)$. Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, which shows that $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ is bounded. Then, the weak compactness of $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ follows from Lemma 4.7 applied for the weakly compact set $\Omega_{1}:=\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ and the weakly closed, bounded set $\Omega_{2}:=\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$.
@@
$3^{\circ}$. We have $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})\subseteq\operatorname{cl}(\operatorname{conv}\,C)=\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}\,S_{C}(||\cdot||)$, where the latter equality is discussed in [29, Rem. 2.5]. Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, hence the set $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ is bounded. Finally, the (weak) compactness of $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ follows from Lemma 4.7 applied for the (weakly) compact set $\Omega_{1}:=\operatorname{cl}\,S_{C}(||\cdot||)$ and the (weakly) closed, bounded set $\Omega_{2}:=\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$.
@@
$4^{\circ}$. Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})\subseteq\operatorname{cl}\,S_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, hence both weakly closed sets $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ and $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ are weakly compact if $\operatorname{\mathbb{B}}_{||\cdot||}$ is weakly compact.
@@
To later prove a Bishop-Phelps type scalarization result (Corollary 5.30) for vector optimization problems in real normed spaces, we need the following proposition.
Proposition 4.11.
@@
Consider a real normed space $(Y,||\cdot||)$ with dimension two or higher, a nontrivial cone $K\subseteq Y$ with the norm-base $B_{K}(||\cdot||)$, a nontrivial cone $A\subseteq Y$ with the norm-base $B_{A}(||\cdot||)$. Suppose that (4.2) and (4.3) with $\psi(\cdot):=||\cdot||$ are valid. Moreover, assume that one of the following statements is valid:
@@
Then, there is $||\cdot||^{\prime}\sim||\cdot||$ such that (4.1) with $\psi(\cdot):=||\cdot||^{\prime}$ is valid.
Proof 4.12.
Suppose that (4.2) and (4.3) (with $\psi(\cdot):=||\cdot||$) are valid. First, note that the following assertions are known to be equivalent (see [14, Lem. 1.25 and 1.26], [29, Lem. 3.4] and [54, Th. 3.2]):
  • $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$ (i.e., (4.3) with $\psi(\cdot):=||\cdot||$ is valid);

### v3

@@ 4.2 Nonlinear Cone Separation Conditions
The key tool to prove our main Bishop-Phelps type scalarization results in vector optimization will be the separation of two (not necessarily convex) cones by using Bishop-Phelps type separating cones / separating (seminorm-linear) functions. In this section we present all the necessary basics for nonlinear cone separation.
@@
$(\operatorname{cl}\,S_{A}^{0}(\psi))\cap(\operatorname{cl}\,S_{-K}(\psi))=\emptyset,$ (10)
$A\cap(\operatorname{cl}(\operatorname{conv}(-K))=\{0\},$ (11)
$0\notin\operatorname{cl}\,S_{-K}(\psi).$ (12)
Remark 13.
We would like to point out that a separation condition for two (not necessarily convex) cones,
$(\operatorname{cl}\,S_{{\rm bd}\,A}^{0}(\psi))\cap(\operatorname{cl}\,S_{-K}(\psi))=\emptyset,$ (13)
which is a similar condition to the one in (10) (more precisely, $A$ is replaced by ${\rm bd}\,A$), was initially studied by Kasimbeyli [8, Def. 4.1, Th. 4.3], [9] and more recently by García-Castaño, Melguizo-Padial and Parzanese [1] in the context of cone separation based on Bishop-Phelps separating cones / functions in real normed spaces (with $\psi$ given by the underlying norm of the space). Moreover, in the same framework, Günther, Khazayel and Tammer [15] studied the separation condition (10) and compared their results with those results (from [8] and [1]) obtained for the separation condition (13). In [14], the study of the separation condition (10) and cone separation based on Bishop–Phelps type separating cones / functions is conducted within the more general framework of real topological-linear spaces.
[…]
Proposition 19.
Assume that (11) and (12) are valid, and one of the following statements is satisfied:
@@
    $\operatorname{cl}\,S_{-K}(\psi)$ =$=\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}$ (14)
@@
    $B_{A}\text{ is compact},$ (15)
    $S_{A}^{0}\text{ is closed},$ (16)
    $A\text{ is closed and }\operatorname{cl}\,B_{A}\text{ is compact},$ (17)
    $A\text{ is closed and }\operatorname{conv}(\operatorname{cl}(B_{A})\cup\{0\})\text{ is closed}.$ (18)
Then, (10) is valid.
Proof.
a) Assuming $A$ is closed and convex, by (11) we get $\operatorname{cl}\,S_{A}^{0}\cap\operatorname{cl}\,S_{-K}\subseteq A\cap\operatorname{cl}(\operatorname{conv}(-K))=\{0\}$, and so by (12) we conclude (10).
@@
b) Obviously, (11) and (12) imply $A\cap\operatorname{cl}\,S_{-K}=\emptyset$. Since there is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ with (14), we get that $x^{*}(x)<\alpha$ for all $x\in A\cap(\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi})$. Note that $B_{A}\cup\{0\}\subseteq A\cap\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}$, and if $A$ is closed, then $\operatorname{cl}\,B_{A}\cup\{0\}\subseteq A\cap\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}$. Consider cases:
@@
Case 1: (15) is valid. By the continuity of $x^{*}$ and the compactness of $B_{A}\cup\{0\}$, we get some $\beta\in(0,\alpha)$ such that $B_{A}\cup\{0\}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\leq\beta\}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$. Hence, $\operatorname{cl}\,S_{A}^{0}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\leq\beta\}$.
Case 2: (16) is valid. Then, $\operatorname{cl}\,S_{A}^{0}=S_{A}^{0}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$.
@@
Case 3: (17) is valid. By the continuity of $x^{*}$ and the compactness of $\operatorname{cl}\,B_{A}\cup\{0\}$ we get some $\beta\in(0,\alpha)$ such that $\operatorname{cl}\,B_{A}\cup\{0\}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\leq\beta\}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$. Thus, $\operatorname{cl}\,S_{A}^{0}\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\leq\beta\}$.
@@
Case 4: (18) is valid. Then, $\operatorname{cl}\,S_{A}^{0}\subseteq\operatorname{cl}(\operatorname{conv}(\operatorname{cl}(B_{A})\cup\{0\}))=\operatorname{conv}(\operatorname{cl}(B_{A})\cup\{0\})\subseteq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)<\alpha\}$.
Recalling that (14) is valid, we conclude (10). ∎
Remark 14.
Note that the equality “$=$” in (14) of Proposition 19 is essential and can not be replaced by $\subseteq$”. In the left-hand side of Figure 1, conditions (11), (12), (14), (15) are valid and the expected conclusion given by (10) holds. In the right-hand side of the Figure 1, the conditions (11), (12), (15) are valid, there is no $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ such that condition (14) is valid, but there is a pair $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ such that
$\operatorname{cl}\,S_{-K}(\psi)\subsetneq\{x\in\operatorname{cl}\,\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}$
holds. In this case, the conclusion given by (10) fails. This example encourages studying equivalent norms in the contexts of cone separation (see the results in this section) and scalarization (see Section 4.4).
${\rm cl}\,S_{A}(||\cdot||_{2})$$A$$A$$-K$${\rm cl}\,S_{-K}(||\cdot||_{2})$$A$$A$$-K$$\{x\in\mathbb{R}^{2}\mid||x||_{2}=1\}$$\{x\in\mathbb{R}^{2}\mid||x||_{1}=1\}$$0$$\{x\in\mathbb{R}^{2}\mid x^{*}(x)=\alpha\}$${\rm cl}\,S_{A}(||\cdot||_{1})$${\rm cl}\,S_{-K}(||\cdot||_{1})$$\{x\in\mathbb{R}^{2}\mid x^{*}(x)=\alpha\}$
Figure 1: Examples given in Remark 14.
Remark 15.
@@
The result in Proposition 19 (regarding assumption b)) is a generalization of a result by Kasimbeyli [9, Lem. 3], where the author studied the Euclidean space $Y=\operatorname{\mathbb{R}}^{n}$, a closed cone $A\subseteq Y$ and a closed convex cone $K\subseteq Y$ with $A\cap(-K)=\{0\}$. Moreover, it is assumed that there is $(x^{*},\alpha)\in Y^{*}\times\operatorname{\mathbb{P}}$ such that (14) is valid (hence $0\notin\operatorname{cl}\,S_{-K}$ is valid). The conclusion in [9, Lem. 3] is given as the separation condition (13) (i.e., (10) with $\operatorname{bd}\,A$ in the role of $A$).
@@
The condition (14) involved in b) of Proposition 19 can be fulfilled by certain Bishop-Phelps type cones, as the following lemma shows.
Lemma 20.
@@
Case 2: Assume that $\psi(d)<1$. Put $\gamma:=x^{*}(d)\,(\geq\alpha)$. By our assumption (19), there is $e\in Y\setminus\{0\}$ with $\psi(e)=1$ and $x^{*}(e)=0$. Then, $x^{*}(d\pm te)=x^{*}(d)\pm tx^{*}(e)=\gamma$ for all $t\geq 0$, and $\psi(d\pm te)\geq|\,\psi(d)-\psi(te)\,|=|\,\psi(d)-|t|\,|\to\infty$ for $t\to\infty$. Taking into account the continuity of $t\mapsto h_{+}(t):=\psi(d+te)$ and $t\mapsto h_{-}(t):=\psi(d-te)$, as well as the conditions $h_{+}(0)=h_{-}(0)=\psi(d)<1$, $\lim_{t\to\infty}\,h_{+}(t)=\lim_{t\to\infty}\,h_{-}(t)=\infty$, by the well-known intermediate value theorem there are $t_{+},t_{-}\in\operatorname{\mathbb{P}}$ with $h_{+}(t_{+})=h_{-}(t_{-})=1$. Since $x^{*}(d+t_{+}e)=x^{*}(d-t_{-}e)=\gamma\geq\alpha$ we infer $d-t_{-}e\in B_{C}$ and $d+t_{+}e\in B_{C}$, hence $d\in\operatorname{conv}(\{d-t_{-}e,\,d+t_{+}e\})\subseteq S_{C}.$
Of course, $S_{C}=D$ and the closedness of $D$ imply $\operatorname{cl}\,S_{C}=S_{C}$. ∎
Remark 16.
Kasimbeyli and Kasimbeyli [51, Th. 4.5] stated a similar result as in Lemma 20 in a reflexive Banach space setting (with $\psi(\cdot):=||\cdot||$). Note that the assumption given in (19), which in this setting means that $Y$ has dimension two or higher, is essential. In the one dimensional space ($Y:=\mathbb{R}$, ||$\cdot$|| ⋅ |), the condition (19) fails and the conclusion of Lemma 20 may fail. To see this, consider $Y:=\mathbb{R}$, $x^{*}:=1$ and $\alpha\in(0,1]$. Then, $C=C(x^{*},\alpha)=\mathbb{R}_{+}$ and $B_{C}(|\cdot|)=\{1\}$ as well as $\operatorname{cl}\,S_{C}(|\cdot|)=\{1\}$ while $\{x\in\operatorname{\mathbb{B}}_{\psi}\mid x^{*}(x)\geq\alpha\}=[\alpha,1]$.
To prove some more specific results in the normed space setting in the upcoming Proposition 23 and Corollary 33 (see Section 4.4), we need additional lemmata.
Lemma 21.
@@
$\Omega_{2}\subseteq[0,t]\cdot\Omega_{1}.$ (20)
@@
Define the function $h:\operatorname{\mathbb{R}}\times Y\to Y$ by $h(s,\omega):=s\cdot\omega$ for all $(s,\omega)\in\operatorname{\mathbb{R}}\times Y$. Since $[0,t]\times\Omega_{1}$ is (weakly) compact in ×$\operatorname{\mathbb{R}}\times Y$, and the function $h$ is (weakly) continuous by the topological vector space properties, we get that $[0,t]\cdot\Omega_{1}$ is (weakly) compact. From (20) we derive that the (weakly) closed subset $\Omega_{2}$ of $[0,t]\cdot\Omega_{1}$ is (weakly) compact as well. ∎
Lemma 22.
@@
Proof.
@@
Consider any $||\cdot||^{\prime}\sim||\cdot||$. $1^{\circ}$ Obviously, $B_{C}(||\cdot||^{\prime})\subseteq C=\operatorname{\mathbb{R}}_{+}\cdot B_{C}(||\cdot||)$. Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $B_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, which shows that $B_{C}(||\cdot||^{\prime})$ is bounded. Moreover, according to [21, Lemma 2.1.43 (iii)], for the bounded and closed norm-base $B_{C}(||\cdot||)$ with $0\notin B_{C}(||\cdot||)$ we have $C=\mathbb{R}_{+}\cdot B_{C}(||\cdot||)$ is closed, hence $B_{C}(||\cdot||^{\prime})$ is closed as well. Then, the compactness of $B_{C}(||\cdot||^{\prime})$ follows from Lemma 21 applied for the compact set $\Omega_{1}:=B_{C}(||\cdot||)$ and the closed, bounded set $\Omega_{2}:=B_{C}(||\cdot||^{\prime})$. Note that $0\notin B_{C}(||\cdot||)=\Omega_{1}$.
@@
$2^{\circ}$ Since $C=\operatorname{\mathbb{R}}_{+}\cdot B_{C}(||\cdot||)\subseteq\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||)\subseteq\operatorname{cl}_{w}\,C$, we have $\operatorname{cl}_{w}\,C=\operatorname{cl}_{w}(\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||))$. Taking into account that $\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ is a weakly closed, bounded set with $0\notin\operatorname{cl}_{w}\,B_{C}(||\cdot||)$, applying [21, Lemma 2.1.43 (iii)] we conclude that $\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ is weakly closed (remark that the weak topology of the normed space $Y$ is Hausdorff). Consequently, $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})\subseteq\operatorname{cl}_{w}\,C=\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}_{w}\,B_{C}(||\cdot||)$. Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, which shows that $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ is bounded. Then, the weak compactness of $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ follows from Lemma 21 applied for the weakly compact set $\Omega_{1}:=\operatorname{cl}_{w}\,B_{C}(||\cdot||)$ and the weakly closed, bounded set $\Omega_{2}:=\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$.
@@
$3^{\circ}$ We have $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})\subseteq\operatorname{cl}(\operatorname{conv}\,C)=\operatorname{\mathbb{R}}_{+}\cdot\operatorname{cl}\,S_{C}(||\cdot||)$, where the latter equality is discussed in [15, Rem. 2.5]. Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, hence the set $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ is bounded. Finally, the (weak) compactness of $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ follows from Lemma 21 applied for the (weakly) compact set $\Omega_{1}:=\operatorname{cl}\,S_{C}(||\cdot||)$ and the (weakly) closed, bounded set $\Omega_{2}:=\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$.
@@
$4^{\circ}$ Since $||\cdot||^{\prime}\sim||\cdot||$, there is $s>0$ such that $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})\subseteq\operatorname{cl}\,S_{C}(||\cdot||^{\prime})\subseteq\operatorname{\mathbb{B}}_{||\cdot||^{\prime}}\subseteq s\cdot\operatorname{\mathbb{B}}_{||\cdot||}$, hence both weakly closed sets $\operatorname{cl}_{w}\,B_{C}(||\cdot||^{\prime})$ and $\operatorname{cl}\,S_{C}(||\cdot||^{\prime})$ are weakly compact if $\operatorname{\mathbb{B}}_{||\cdot||}$ is weakly compact. ∎
@@
To later prove a Bishop-Phelps type scalarization result (Corollary 33) for vector optimization problems in real normed spaces, we need the following proposition.
Proposition 23.
@@
Consider a real normed space $(Y,||\cdot||)$ with dimension two or higher, a nontrivial cone $K\subseteq Y$ with the norm-base $B_{K}(||\cdot||)$, a nontrivial cone $A\subseteq Y$ with the norm-base $B_{A}(||\cdot||)$. Suppose that (11) and (12) with $\psi(\cdot):=||\cdot||$ are valid. Moreover, assume that one of the following statements is valid:
@@
Then, there is $||\cdot||^{\prime}\sim||\cdot||$ such that (10) with $\psi(\cdot):=||\cdot||^{\prime}$ is valid.
Proof.
Suppose that (11) and (12) (with $\psi(\cdot):=||\cdot||$) are valid. First, note that the following assertions are known to be equivalent (see [54, Lem. 1.25 and 1.26], [15, Lem. 3.4] and [55, Th. 3.2]):
  • $0\notin\operatorname{cl}\,S_{-K}(||\cdot||)$ (i.e., (12) with $\psi(\cdot):=||\cdot||$ is valid);